# ALMA FACTS III. High-Resolution CO(2-1)/CO(1-0) Maps of Twelve Nearby Galaxies

### v1

@@ 7 Summary
From the radial distributions of , the barred galaxies follow a general trend: is high in the central 1-2 kpc regions 1.0, low along the bar offset ridges 0.6, increases at the bar ends 0.7, and then declines in the rest of the disk with fluctuations along the spiral arms. The radial distributions of the unbarred galaxies fluctuate in and are similar to the outer parts (spiral arms) of the barred galaxies.
High in the central 1-2 kpc regions is always associated with the central gas condensations in the barred galaxies. NGC4826 is the only unbarred spiral galaxy in our sample that shows the central condensation and high . Low along the bar offset ridges suggest that the gas density and/or temperature are lower there, which may explain the reduced star formation efficiency along bar suggested by previous studies. Bar ends show high , indicating increased density and/or temperature. In spiral arms, fluctuates. A part of this fluctuation is due to high-ratio gas around HII regions and thus due to star-formation activity, while an impact of galactic dynamics is another possible factor yet to be quantified.

### v2

@@ VII Summary
From the radial distributions of , the barred galaxies follow a general trend: is high in the central 1-2 kpc regions 1.0, low along the bar offset ridges 0.6, increases at the bar ends often 0.7, and then declines in the rest of the disk with fluctuations along the spiral arms. The radial distributions of the unbarred galaxies fluctuate in and are similar to the outer parts (spiral arms) of the barred galaxies.
High in the central 1-2 kpc regions is always associated with the central gas condensations in the barred galaxies. NGC 4826 is the only unbarred spiral galaxy in our sample that shows the central condensation and high . Low along the bar offset ridges suggest that the gas density and/or temperature are lower there (Koda et al., 2012), which may explain the reduced star formation efficiency along bars as suggested by previous studies. Bar ends show high , indicating increased density and/or temperature. In spiral arms, fluctuates, presumably due to the impacts of star-formation activity or galactic dynamics.